# CushionCatch: A Compliant Catching Mechanism for Mobile Manipulators via Combined Optimization and Learning

### v3

@@ Abstract
Catching flying objects with a cushioning process is a skill commonly performed by humans, yet it remains a significant challenge for robots. In this paper, we present a framework that combines optimization and learning to achieve compliant catching on mobile manipulators (CCMM). First, we propose a high-level capture planner for mobile manipulators (MM) that calculates the optimal capture point and joint configuration. Next, the pre-catching (PRC) planner ensures the robot reaches the target joint configuration as quickly as possible. To learn compliant catching strategies, we propose a network that leverages the strengths of LSTM for capturing temporal dependencies and positional encoding for spatial context (P-LSTM). This network is designed to effectively learn compliant strategies from human demonstrations. Following this, the post-catching (POC) planner tracks the compliant sequence output by the P-LSTM while avoiding potential collisions due to structural differences between humans and robots. We validate the CCMM framework through both simulated and real-world ball-catching scenarios, achieving a success rate of 98.70% in simulation, 92.59% in real-world tests, and a 28.7% reduction in impact torques. The open source code will be released for the reference of the community1.
@@ I INTRODUCTION
3. Simulation and physical experiments were conducted on a mobile manipulator to validate the effectiveness of our framework and the role of compliance. A catching success rates of 98.70% in simulation, 92.59% in physical experiments and a 28.7% reduction in impact torque were achieved. An open source code of the proposed method will be released1.
@@ III-A System Model
[…]
@@ IV-B High-Level Capture Planner
The high-level planner outputs an optimal catching configuration $\bm{q}_{ca}\in\mathbb{R}^{n}$ and catch time $t_{ca}\in\mathbb{R}$ that enables the container to intercept the ball’s trajectory at time $t_{ca}$. From current state $\bm{q}_{0}$, catching a moving ball requires the container and ball to be aligned in position and parallel in the direction vector. For this purpose, the following QP problem is formulated:
@@
${}_{c}^{w}\bm{P}\in\mathbb{R}^{3}$ represents the position of the container in the world frame. Equation (5b) ensures that the position of the container is the same as the ball’s position at $\bm{q}_{ca}$ ${}_{c}^{w}\bm{z}(\bm{q}_{ca})\in\mathbb{R}^{3}$ indicates the container’s z-axis vector expressed in the world frame. Through (5c), the direction of the container is opposite to the velocity of the ball at $t_{ca}$, that is, parallel.
@@ IV-C PRC Planner
where Poly5() represents a Quintic Polynomial Planning that maps the time $t$(ranging from 0 to $T_{prc}$) to the joint position based on the inputs $T_{prc},\bm{q}_{0},\bm{q}_{ca}$.
@@ IV-D Compliant Sequence Learning
where |||| means concatenation horizontally and $k$ is the embedding dimension.
@@ IV-E POC Planner
Figure 7: Snapshots of some catching experiments in the simulation where robot opacity increased with time. It shows the ball approaching the robot from the front, right, left, and back, respectively.
@@ V EXPERIMENTS
We evaluate our framework through experiments conducted both in simulation and on a physical mobile manipulator. We use the following value for the parameters of our framework: $k_{ad}=0.0295$ in (3); $\alpha=2$ in (5a); $\beta=0.5$ in (5d); ${\mathit{\Lambda}}_{b}=5,{\mathit{\Lambda}}_{m}=1$ in (6); $\lambda=1.5$ in (8); $\gamma_{z}=0.1$ in (12c); $\gamma_{xy}=0.1$ in (12d). We use ”SLSQP” method to solve nonlinear optimization problems (5a) and (12a). In P-LSTM, we set the hidden size to 64 and a fully connected layer block from 64 to 6. During training, we use mean squared error (MSE) between the predicted output and the ground truth. The learning rate is set to 0.001, and the max output sequence length is capped at 16.
@@ V-B Network Comparison Experiment
As shown in TABLE I and Fig. 6, transformer-based methods struggle to converge and validate effectively due to their high parameter demands and the short input sequences. The diffusion policy also fails to train effectively due to the limited data size. Additionally, adding PE to both LSTM and transformer models reduces the error, with almost no impact on inference speed in test experiments.”
@@ V-C Simulation Scenarios
We use the Swift and Rtb Python libraries [34] as our simulated environments. Initially, we generated 6,000 ball samples $[p_{x},p_{y},p_{z},\dot{p}_{x},\dot{p}_{y},\dot{p}_{z}]$, restricting their positions within a cylindrical space of a specified height. We ensured that their trajectories, for $t>0.5$, would pass through the robot’s capture range. Then ball parameters are generated randomly by procedure. Additionally, we performed ablation studies on the collision avoidance constraints in (12c) and (12d) during the POC phase (2,000 pieces of data for each test). The statistical results are presented in TABLE II.
@@
Fig. 7 shows the robot motions of some experiments where the ball comes from different directions. TABLE II demonstrates the results of 6000 random catching tasks, showing that the CCMM framework achieves a 98.70% success rate for catching. In practice, constraints (12c) and (12d) do not affect the PRC planner, resulting in similar ”Not Catch” rates. However, incorporating collision avoidance constraints notably reduces the incidence of collisions. Without (12c) or (12d), the probability of collisions in their respective dimensions increases substantially.
@@ V-D Real-World Scenarios
We conducted a total of 100 hand-throwing experiments for compliant catching. We assess the effectiveness of the hand-thrown ball based on the closest distance between the container’s initial position and the ball’s trajectory. Due to inaccuracies in hand throwing, 19 attempts were deemed invalid. Out of the remaining 81 valid throws, the robot successfully captured the ball 75 times, resulting in a success rate of 92.59%, without any collision. The failure occurred because noise in the capture system caused a discrepancy between the filter’s estimated state and the actual state, leading to deviations in the predicted trajectory. It is worth mentioning that, due to the limitation and some obstacles of the laboratory room (some areas wound do not capture the ball well), we did not test in all directions like the simulation, but we still tried to throw to different directions of the robot (left, right, top and middle). In all experiments, TABLE III presents the error between the container’s capture point and the actual ball position ”cap. Error”, along with the capture time ”cap. Time” and the total calculation time ”cal. Time” for the Hive-Level, the PRC, and the Network inference phase. TABLE IV provides collision avoidance data, including the minimum distances between the container and the ground “Min Ground Dis.”, as well as the container and the mobile base ”Min Base Dis.”, during the POC phase. We successfully achieved a high precision capture and collision avoidance.
@@
Due to different hardware in each joint, every torque was normalized according to the maximum allowed torque at that joint, thus, ${\tau^{\prime}}_{i}=\tau_{i}/\tau_{i}^{\max},i\in[2,8]$. We calculate the maximum peak torque $max(\tau^{\prime}_{i})$ in all joints during the experiment, as the ”Max Peak Torque” in TABLE VI. The p-value for the ”Max Peak Torque” between the two groups is significantly lower than the 0.05 significance level, indicating a notable difference between the groups and confirming the feasibility of our compliant strategy. Using the mean value calculation, the quantitative attenuation ratio is $(0.87-0.62)/0.87=28.7\%$. Since there is currently no research specifically on compliant catching for mobile manipulators, we referred to the compliant ratio data for a fixed manipulator catching a vertically falling object [7], which is 28%. This result is similar to ours. However, our study is conducted in a more complex catching scenario and with a mobile manipulator. For one test in compliant catching, Figure 9(a) shows the expected output of the high-level capture planner, along with the changes in joint angles during the catching process. At the same time, Figure 9(b) presents the actual predicted trajectory of the ball and the container’s trajectory during the compliant catching experiment. Additionally, Fig. 10 illustrates the normalized joint torques.

### v4

@@ Abstract
Catching flying objects with a cushioning process is a skill commonly performed by humans, yet it remains a significant challenge for robots. In this paper, we present a framework that combines optimization and learning to achieve compliant catching on mobile manipulators (CCMM). First, we propose a high-level capture planner for mobile manipulators (MM) that calculates the optimal capture point and joint configuration. Next, the pre-catching (PRC) planner ensures the robot reaches the target joint configuration as quickly as possible. To learn compliant catching strategies, we propose a network that leverages the strengths of LSTM for capturing temporal dependencies and positional encoding for spatial context (P-LSTM). This network is designed to effectively learn compliant strategies from human demonstrations. Following this, the post-catching (POC) planner tracks the compliant sequence output by the P-LSTM while avoiding potential collisions due to structural differences between humans and robots. We validate the CCMM framework through both simulated and real-world ball-catching scenarios, achieving a success rate of 98.70% in simulation, 92.59% in real-world tests, and a 28.7% reduction in impact torques. The open source code has be released for the reference of the community1.
@@ I INTRODUCTION
3. Simulation and physical experiments were conducted on a mobile manipulator to validate the effectiveness of our framework and the role of compliance. A catching success rates of 98.70% in simulation, 92.59% in physical experiments and a 28.7% reduction in impact torque were achieved.
@@ III-A System Model
where $\bm{p}\in\mathbb{R}^{3}$ is the ball’s position in the world frame, $\bm{g}=[0,0,g]^{\text{T}}$ is the gravitational acceleration, $k_{ad}$ is the aerodynamic drag coefficient and $\|\cdot\|$ is the Euclidean norm.
@@ IV-B High-Level Capture Planner
The high-level planner outputs an optimal catching configuration $\bm{q}_{ca}\in\mathbb{R}^{n}$ and catch time $t_{ca}\in\mathbb{R}$ that enables the container to intercept the ball’s trajectory at time $t_{ca}$. From current state $\bm{q}_{0}$, catching a moving ball requires the container and ball to be aligned in position and parallel in the direction vector. For this purpose, the following optimization problem is formulated:
@@
$_{c}^{w}\bm{P}\in\mathbb{R}^{3}$ represents the position of the container in the world frame. Equation (5b) ensures that the position of the container is the same as the ball’s position at $\bm{q}_{ca}$ . $_{c}^{w}\bm{z}(\bm{q}_{ca})\in\mathbb{R}^{3}$ indicates the container’s z-axis vector expressed in the world frame. Through (5c), the direction of the container is opposite to the velocity of the ball at $t_{ca}$, that is, parallel.
@@ IV-C PRC Planner
where Poly5() represents a Quintic Polynomial Planning that maps the time tt(ranging from 0 to $T_{prc}$) to the joint position based on the inputs $T_{prc},\bm{q}_{0},\bm{q}_{ca}$.
@@ IV-D Compliant Sequence Learning
where $|$ means concatenation horizontally and $k$ is the embedding dimension.
@@ IV-E POC Planner
Figure 7: Snapshots from simulation experiments show the ball approaching the robot from various directions, with the robot’s opacity gradually increasing over time.
@@ V EXPERIMENTS
We evaluate our framework through experiments conducted both in simulation and on a physical mobile manipulator. We use the following value for the parameters of our framework: $k_{ad}=0.0295$ in (3); $\alpha=2$ in (5a); $\beta=0.5$ in (5d); ${\mathit{\Lambda}}_{b}=5,{\mathit{\Lambda}}_{m}=1$ in (6); $\lambda=1.5$ in (8); $\gamma_{z}=0.1$ in (12c); $\gamma_{xy}=0.1$ in (12d). We use “SLSQP” method to solve nonlinear optimization problems (5a) and (12a). In P-LSTM, we set the hidden size to 64 and a fully connected layer block from 64 to 6. During training, we use mean squared error (MSE) between the predicted output and the ground truth. The learning rate is set to 0.001, and the max output sequence length is capped at 16.
@@ V-B Network Comparison Experiment
As shown in TABLE I and Fig. 6, transformer-based methods struggle to converge and validate effectively due to their high parameter demands and the short input sequences. The diffusion policy also fails to train effectively due to the limited data size. Additionally, adding PE to both LSTM and transformer models reduces the error, with almost no impact on inference speed in test experiments.
@@ V-C Simulation Scenarios
We use the Swift and Rtb Python libraries [34] as our simulated environments. Initially, we generated 6,000 ball samples $p_{x},p_{y},p_{z},\dot{p}_{x},\dot{p}_{y},\dot{p}_{z}]$, restricting their positions within a cylindrical space of a specified height. We ensured that their trajectories, for $t>0.5$, would pass through the robot’s capture range. Then ball parameters are generated randomly by procedure. Additionally, we performed ablation studies on the collision avoidance constraints in (12c) and (12d) during the POC phase (2,000 pieces of data for each test). The statistical results are presented in TABLE II.
@@
Fig. 7 shows the robot motions of some experiments where the ball comes from different directions. TABLE II demonstrates the results of 6000 random catching tasks, showing that the CCMM framework achieves a 98.70% success rate for catching. In practice, constraints (12c) and (12d) do not affect the PRC planner, resulting in similar “Not Catch” rates. However, incorporating collision avoidance constraints notably reduces the incidence of collisions. Without (12c) or (12d), the probability of collisions in their respective dimensions increases substantially.
@@ V-D Real-World Scenarios
We conducted a total of 100 hand-throwing experiments for compliant catching. We assess the effectiveness of the hand-thrown ball based on the closest distance between the container’s initial position and the ball’s trajectory. Due to inaccuracies in hand throwing, 19 attempts were deemed invalid. Out of the remaining 81 valid throws, the robot successfully captured the ball 75 times, resulting in a success rate of 92.59%, without any collision. The failure occurred because noise in the capture system caused a discrepancy between the filter’s estimated state and the actual state, leading to deviations in the predicted trajectory. It is worth mentioning that, due to the limitation and some obstacles of the laboratory room (some areas wound do not capture the ball well), we did not test in all directions like the simulation, but we still tried to throw to different directions of the robot (left, right, top and middle). In all experiments, TABLE III presents the error between the container’s capture point and the actual ball position “cap. Error”, along with the capture time “cap. Time” and the total calculation time “cal. Time” for the Hive-Level, the PRC, and the Network inference phase. TABLE IV provides collision avoidance data, including the minimum distances between the container and the ground “Min Ground Dis.”, as well as the container and the mobile base “Min Base Dis.”, during the POC phase. We successfully achieved a high precision capture and collision avoidance.
@@
Due to different hardware in each joint, every torque was normalized according to the maximum allowed torque at that joint, thus, ${\tau^{\prime}}_{i}=\tau_{i}/\tau_{i}^{\max},i\in[2,8]$. We calculate the maximum peak torque $max(\tau^{\prime}_{i})$ in all joints during the experiment, as the “Max Peak Torque” in TABLE VI. The p-value for the “Max Peak Torque” between the two groups is significantly lower than the 0.05 significance level, indicating a notable difference between the groups and confirming the feasibility of our compliant strategy. Using the mean value calculation, the quantitative attenuation ratio is $(0.87-0.62)/0.87=28.7\%$. Since there is currently no research specifically on compliant catching for mobile manipulators, we referred to the compliant ratio data for a fixed manipulator catching a vertically falling object [7], which is 28%. This result is similar to ours. However, our study is conducted in a more complex catching scenario and with a mobile manipulator. For one test in compliant catching, Figure 9(a) shows the expected output of the high-level capture planner, along with the changes in joint angles during the catching process. At the same time, Figure 9(b) presents the actual predicted trajectory of the ball and the container’s trajectory during the compliant catching experiment. Additionally, Fig. 10 illustrates the normalized joint torques.